OpenStudy (mony01):

The position of a particle is given by the expression x(t) = 4.00 cos(3.00pi(t) + pi), where x is in meters and t is in seconds. Determine the frequency.

OpenStudy (anonymous):

now remember the definition of frequency is 1 over the period

OpenStudy (anonymous):

use the standard definition of a cosine function to help you determine the period http://www.regentsprep.org/regents/math/algtrig/att7/sinusoidal.htm

OpenStudy (anonymous):

with this you place 1 over the period and you have your frequency

OpenStudy (mony01):

is the period 2pi/pi?

OpenStudy (mony01):

@recon14193

OpenStudy (anonymous):

standard cosine equation would be\[Acos(b(x-c))+d\]

OpenStudy (anonymous):

period would be\[\frac{ 2\pi }{ b }\]

OpenStudy (anonymous):

here we have\[4\cos(3\pi t+\pi)\]

OpenStudy (anonymous):

if I understand you correctly that is the equation and so b should be *badum tss* in this instance 2pi/3pi

OpenStudy (anonymous):

or excuse me b should be 3pi so our period is 2pi/3pi